OpenStudy (theslytherinhelper):

How does √6 = 6 (1/2)?

OpenStudy (anonymous):

sqrt 6= 6^(1/2) not 6(1/2)

OpenStudy (theslytherinhelper):

Yeah, I forgot to put the '^', but I don't understand how :/

OpenStudy (anonymous):

This is because of the squareroot sign

OpenStudy (ryan123345):

\[6^{1/2}=\sqrt{6}\]

OpenStudy (anonymous):

so for eg:- if you 64 you know 8^2 =64 the sqroot sign is just inverse of square so inverse of square is 1/2 that is why sqrt 6 = 6^(1/2)

OpenStudy (theslytherinhelper):

Ohhhhhh, so basically, in order to get rid of the square root sign in an equation, I'd have to make the exponent 1/2?

OpenStudy (ryan123345):

No. In order to get rid of the square root sign you have to square it (^2)

OpenStudy (ryan123345):